Hector Elizondo took out a 65-day loan of $7500 at an annual interest rate of 5.5%. Find the simple interest due on the loan. (Round your answer to the nearest cent.)
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

Interest = P * R *T
convert days to years years = 65/365
Interest = 7500*5.5% * (65/365)
= $73.5

Answer by ikleyn(53944)

.
Hector Elizondo took out a 65-day loan of $7500 at an annual interest rate of 5.5%.
Find the simple interest due on the loan. (Round your answer to the nearest cent.)
~~~~~~~~~~~~~~~~~~~~~~~


        The assignment requires the answer to the nearest cent.
        @mananth failed this request.
        See my correct calculations below.


Interest = P * R *T

convert days to years = 65/365

Interest = 7500 * 0.055 * (65/365)

= $73.46 to the nearest cent.         ANSWER


So, @mananth either does not read the assignment to the end, or his code makes rounding
based on its own rules that @mananth does not control to the end - and then @mananth does not read/look at the output of his code.